# Exact Exploratory Bi-factor Analysis: A Constraint-based Optimisation Approach

### v1

@@ 2.2 Proposed ALM
Following the previous discussion, we see that we can perform exploratory bi-factor analysis by solving the optimisation problem with some equality constraints and the constraint that $\Phi$ is a correlation matrix. To deal with the constraints in $\Phi$, we consider the following reparameterisation that has been considered in Alfonzetti et al., (2024), which is also similar to the implementation in the state-of-the-art statistical software Stan (Stan Development Team,, 2022):
$\Phi=\left(\begin{bmatrix}1&\mathbf{0}^{\top}\\ \mathbf{0}&U^{T}\end{bmatrix}\right)\left(\begin{bmatrix}1&\mathbf{0}^{\top}\\ \mathbf{0}&U\end{bmatrix}\right),$ (4)
where $U$ is defined recursively by
[…]
With slight abuse of notation, we reexpress the covariance matrix as $\Phi(\bm{\gamma})$. In addition, we use $\bm{\psi}=(\psi_{1},...,\psi_{J})^{\top}$ to denote the vector of diagonal entries of $\Psi$ and reexpress the residual covariance matrix as $\Psi(\bm{\psi})$. Thus, the optimisation problem (2) is now simplified as
@@
$\min_{\Lambda,\bm{\gamma},\bm{\psi}}~{}~{}~{}l(\Lambda\Phi(\bm{\gamma})\Lambda^{\top}+\Psi(\bm{\psi});S)+\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}\beta_{jkk^{\prime}}^{(t)}\lambda_{jk}\lambda_{jk^{\prime}}\right)+c^{(t)}\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}(\lambda_{jk}\lambda_{jk^{\prime}})^{2}\right),$ (7)
where $c^{(t)}$ and $\beta_{jkk^{\prime}}^{(t)}$s are auxiliary coefficients of the ALM determined by the result of the $(t-1)$th optimisation. Details of the ALM are given in Algorithm 1 below, where the function $h$ returns the second-largest value of a vector.
@@
$=$ $\arg\min_{\Lambda,\bm{\gamma},\bm{\psi}}~{}~{}~{}l(\Lambda\Phi(\bm{\gamma})\Lambda^{\top}+\Psi(\bm{\psi});S)+\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}\beta_{jkk^{\prime}}^{(t)}\lambda_{jk}\lambda_{jk^{\prime}}\right)+c^{(t)}\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}(\lambda_{jk}\lambda_{jk^{\prime}})^{2}\right).$
4:     if $d\left((\Lambda^{(t+1)},\bm{\gamma}^{(t+1)},\bm{\psi}^{(t+1)}),(\Lambda^{(t)},\bm{\gamma}^{(t)},\bm{\psi}^{(t)})\right)<\epsilon$ then
5:         return $\Lambda^{(t+1)},\bm{\gamma}^{(t+1)},\bm{\psi}^{(t+1)}$.
6:     end if
7:     $\beta_{jkk^{\prime}}^{(t+1)}=\beta_{jkk^{\prime}}^{(t)}+c^{(t)}\lambda^{(t)}_{jk}\lambda^{(t)}_{jk^{\prime}}$;
8:     if $\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}(\lambda_{jk}^{(t)}\lambda_{jk^{\prime}}^{(t)})^{2}\right)^{1/2}>c_{\theta}\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}(\lambda_{jk}^{(t-1)}\lambda_{jk^{\prime}}^{(t-1)})^{2}\right)^{1/2}$ then
9:         $c^{(t+1)}=c_{\sigma}c^{(t)}$;
10:     else
11:         $c^{(t+1)}=c^{(t)}$;
12:     end if
13:     if $\left(\|\Lambda^{(t)}-\Lambda^{(t-1)}\|_{F}^{2}+\|\bm{\gamma}^{(t)}-\bm{\gamma}^{(t-1)}\|^{2}+\|\bm{\psi}^{(t)}-\bm{\psi}^{(t-1)}\|^{2}\right)^{1/2}/\sqrt{J(G+2)+G(G-1)/2}<\delta_{1},$
14:and $\max_{j\in\{1,...,J\}}h(|\lambda_{j2}^{(t)}|,...,|\lambda_{j,G+1}^{(t)}|)<\delta_{2}$ then
15:         Break
16:     end if
17:end while
18:$\Lambda^{(t)},\bm{\gamma}^{(t)},\bm{\psi}^{(t)}$.
The updating rule of $\beta^{(t)}_{jkk^{\prime}}$ and $c^{(t)}$ follows equation (1) and (47) in Chapter 2.2 of Bertsekas, (2014) and the convergence of Algorithm 1 is guaranteed by Proposition 2.7 of Bertsekas, (2014). We follow the recommended choices of $c_{\theta}=0.25$ and $c_{\sigma}=10$ in Bertsekas, (2014).
@@
When both criteria are below their pre-specified thresholds, $\delta_{1}$ and $\delta_{2}$, respectively, we stop the algorithm. Let $T$ be the last iteration number. Then the estimated bi-factor model structure is given by
@@
The optimisation problem 7 is non-convex and can get stuck in a local minimum. Thus, we recommend running the proposed algorithm multiple times with random starting points and choosing the solution with the smallest objective function value. The algorithm can also suffer from slow convergence, especially when the penalty term becomes large. When the algorithm does not converge within $T_{\max}$ iterations, we suggest using the estimated parameters at the $T_{\max}$th iteration as the initial parameters and restarting the optimisation until a good proportion of them converge. In the simulation study in Section 3 below, the estimated parameters obtained using 50 random starting points are close to the global minimum in most cases in the simulation study. For the real data example in Section 4, 200 random starting points are used to ensure a reliable result. We set $T_{\max}=1000$ in all of our numerical studies.
@@ 3.1 Study I
In this study, we compare the proposed method with the oblique bi-factor rotation in Jennrich and Bentler, (2012) regarding their performance in the estimation accuracy of parameters and the recovery of the bi-factor structure. The oblique bi-factor rotation method first estimates the loading matrix $\widehat{\Lambda}$ under the exploratory factor analysis setting by the optimisation problem
@@
We restrict $\lambda_{ij}=0$ for $i=2,\ldots,G$ and $j=i+1,\ldots,G+1$ to avoid the rotational indeterminacy of $\widehat{\Lambda}$, as suggested in Anderson and Rubin, (1956). Then the rotated solutions $\widehat{\Lambda}^{oblq}$ and $\widehat{\Phi}^{oblq}$ are obtained by finding a rotation matrix that solves the optimisation problem for oblique bi-factor rotation (Jennrich and Bentler,, 2012). The implementation in the R package GPArotation (Bernaards and Jennrich,, 2005) is used for solving this optimisation problem, which is based on a gradient projection algorithm. The optimisation problem for rotation is also nonconvex and thus may converge to local solutions. For a fair comparison, we also use 50 random starting points for the initial rotation matrix, as the same number of random starting points are used when running Algorithm 1.
We first examine the accuracy in estimating the loading matrix. We calculate the mean squared error (MSE) for $\widehat{\Lambda}$, after adjusting for the label and sign indeterminacy as considered in Theorem 1 and further discussed in Remark 1. More specifically, let $\mathcal{P}$ and $\mathcal{D}$ be the sets of permutation and sign flip matrices, respectively, as defined in Theorem 1. We define the MSE for $\Lambda$ as
$\text{MSE}_{\Lambda}=\min_{P\in\mathcal{P},D\in\mathcal{D}}\|\widehat{\Lambda}-\Lambda^{*}PD\|_{F}^{2}/(JK).$
This MSE can be applied to the loading matrix estimates from both methods.
@@
for $g=1,\ldots,G$ and some hard thresholding parameter $\delta>0$. We note that ^$\widehat{\mathcal{B}}_{g}^{oblq}$s may not yield an exact bi-factor structure in the sense that the current construction does not ensure them to be mutually exclusive.
Let $\{\mathcal{B}_{g}^{*},g=1,\ldots,G\}$ be the true non-overlapping group partitions of the $J$ variables, and let $\{\mathcal{B}_{g},g=1,\ldots,G\}$ be their estimates, either from the proposed method or the rotation method. As the group factors can only be recovered up to label swapping, as Theorem 1 suggests, we measure the matching between the true and estimated structure up to a permutation of the factor labels. Specifically, the following two evaluation criteria are considered:
@@
  • Average Correctness Criterion (ACC): $\max_{\sigma\in\widetilde{\mathcal{P}}}\sum_{g=1}^{G}(|\mathcal{B}_{g}^{*}\cap\mathcal{B}_{\sigma(g)}|+|\mathcal{B}_{\sigma(g)}^{C}\cap\mathcal{B}_{g}^{*C}|)/(JG)$, which is the average of the number of the bi-factor locations that are correctly learned. Here for any set $\mathcal{B}$, $\mathcal{B}^{C}=\{1,\ldots,J\}\setminus\mathcal{B}$ is the complement of set $\mathcal{B}$.
To examine the recovery of the bi-factor structure, for $(J,G)\in\{(15,3),(30,5)\}$, $N\in\{500,2000\}$ and $\delta\in\{0.2,0.4,0.6,0.8,1\}$, we consider 8 simulation settings. In the data generation model, $\mathcal{B}_{g}^{*}=\{g,g+G,\ldots,g+G(J/G-1)\}$ for $g=1,\ldots,G$, $\Psi^{*}=\mathbb{I}_{J\times J}$, $\Phi^{*}=\Phi^{*}(\bm{\gamma}^{*})$ follows the reparameterization in Section 2.2 with $\bm{\gamma}^{*}$ following i.i.d $N(0,1/4)$ and $\Lambda^{*}$ follows
@@
for $j=1,\ldots,J$ and $k=1,\ldots,G+1$. Here $u_{jk}$ follows i.i.d Uniform(0,1) distribution and $x_{jk}$ follows i.i.d standard normal distribution.
For each setting, we first generate $\Lambda^{*}$ and $\Phi^{*}$ once and use them to generate 100 datasets. The result on the estimation of the loading matrix is shown in Figure 1, and the results of the recovery of the bi-factor structure are shown in Tables 1 and 2. It shows that the proposed method outperforms the rotation method in terms of both criteria. Specifically, according to the EMC criterion, the proposed method has a good chance to completely recover the bi-factor structure, especially when $G$ is small and $N$ is large, while the rotation method with hard thresholding has a much lower chance of achieving so. Regarding the ACC criterion, it suggests that both methods can identify the group factors reasonably well, while the proposed method is more accurate.
Figure 1: Simulation results of the MSE for $\widehat{\Lambda}$ of ALM and exploratory bi-factor rotation.
Table 1: Simulation results of the EMC of ALM and exploratory bi-factor rotation.
$(J,G)$ $N$ ALM $\delta=0.2$ $\delta=0.4$ $\delta=0.6$ $\delta=0.8$ $\delta=1$
(15,3) 500 0.9 0.36 0.00 0.00 0.00 0.00
2000 1.00 0.62 0.00 0.00 0.00 0.00
(30,5) 500 0.70 0.00 0.00 0.00 0.00 0.00
2000 0.96 0.00 0.00 0.00 0.00 0.00
Table 2: Simulation results of the ACC of ALM and exploratory bi-factor rotation.
$(J,G)$ $N$ ALM $\delta=0.2$ $\delta=0.4$ $\delta=0.6$ $\delta=0.8$ $\delta=1$
(15,3) 500 0.99 0.98 0.95 0.93 0.90 0.87
2000 1.00 0.99 0.96 0.93 0.91 0.89
(30,5) 500 0.99 0.84 0.95 0.97 0.95 0.94
2000 0.99 0.89 0.97 0.97 0.95 0.94

### v2

@@ 2.2 Proposed ALM
Following the previous discussion, we see that we can perform exploratory bi-factor analysis by solving the optimisation problem with some equality constraints and the constraint that $\Phi$ is a correlation matrix. To deal with the constraints in $\Phi$, we consider a reparameterisation of $\Phi$ based on a Cholesky decomposition, where the explicit form of the reparametrisation is given in Appendix A. With slight abuse of notation, we reexpress the covariance matrix as $\Phi(\bm{\gamma})$, where $\bm{\gamma}$ is a $G(G-1)/2$ dimensional unconstrained parameter vector. In addition, we use $\bm{\psi}=(\psi_{1},...,\psi_{J})^{\top}$ to denote the vector of diagonal entries of $\Psi$ and reexpress the residual covariance matrix as $\Psi(\bm{\psi})$. Thus, the optimisation problem (2) is now simplified as
@@
$\min_{\Lambda,\bm{\gamma},\bm{\psi}}~{}~{}~{}l(\Lambda\Phi(\bm{\gamma})\Lambda^{\top}+\Psi(\bm{\psi});S)+\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}\beta_{jkk^{\prime}}^{(t-1)}\lambda_{jk}\lambda_{jk^{\prime}}\right)+\frac{1}{2}c^{(t-1)}\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}(\lambda_{jk}\lambda_{jk^{\prime}})^{2}\right),$ (5)
where $c^{(t-1)}>0$ and $\beta_{jkk^{\prime}}^{(t-1)}$s are auxiliary coefficients of the ALM determined by the initial values when $t=1$ and the previous optimisation when $t\geq 2$. Details of the ALM are given in Algorithm 1 below, in which the function $h$ returns the second-largest value of a vector.
The ALM can be seen as a penalty method for solving constrained optimization problems. It replaces a constrained optimization problem with a series of unconstrained problems. It adds a penalty term, i.e., the third term in (5), to the objective to enforce the constraints. The tuning parameter $c^{(t-1)}$ can be seen as the weight of the penalty term. In fact, as $c^{(t)}$ goes to infinity while $\beta_{jkk^{\prime}}^{(t)}$s remain bounded, the solution has to converge to one satisfying the equality constraints in (4), as otherwise, the objective function value in (5) will diverge to infinity. However, the ALM is not purely a penalty method in the sense that it also adds the second term in (5) to mimic a Lagrange multiplier (see, e.g., Chapter 12, Nocedal and Wright,, 1999), for which $\beta_{jkk^{\prime}}^{(t-1)}$s are the weights. An advantage of the ALM is that, with the inclusion of the Lagrangian term (i.e., the second term), the method is guaranteed to converge to a local solution satisfying the equality constraints without requiring $c^{(t)}$ to go to infinity. This is important, as when $c^{(t)}$ is very large, the optimisation problem (5) becomes ill-conditioned and thus hard to solve.
@@
$=$ $\arg\min_{\Lambda,\bm{\gamma},\bm{\psi}}~{}~{}~{}l(\Lambda\Phi(\bm{\gamma})\Lambda^{\top}+\Psi(\bm{\psi});S)+\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}\beta_{jkk^{\prime}}^{(t-1)}\lambda_{jk}\lambda_{jk^{\prime}}\right)+\frac{1}{2}c^{(t-1)}\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}(\lambda_{jk}\lambda_{jk^{\prime}})^{2}\right).$
4:     Update $\beta_{jkk^{\prime}}^{(t)}$ and $c^{(t)}$ according to equations (6) and (7)
$\beta_{jkk^{\prime}}^{(t)}=\beta_{jkk^{\prime}}^{(t-1)}+c^{(t-1)}\lambda^{(t)}_{jk}\lambda^{(t)}_{jk^{\prime}},$ (6)
and
$c^{(t)}=\left\{\begin{array}[]{l}c_{\sigma}c^{(t-1)}\mbox{~{}~{}if~{}~{}}\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}(\lambda_{jk}^{(t)}\lambda_{jk^{\prime}}^{(t)})^{2}\right)^{1/2}>c_{\theta}\left(\sum_{j=1}^{J}\sum_{k=2}^{G}\sum_{k^{\prime}=k+1}^{G+1}(\lambda_{jk}^{(t-1)}\lambda_{jk^{\prime}}^{(t-1)})^{2}\right)^{1/2};\\ c^{(t-1)}\mbox{~{}otherwise}.\end{array}\right.$ (7)
5:     if $\left(\|\Lambda^{(t)}-\Lambda^{(t-1)}\|_{F}^{2}+\|\bm{\gamma}^{(t)}-\bm{\gamma}^{(t-1)}\|^{2}+\|\bm{\psi}^{(t)}-\bm{\psi}^{(t-1)}\|^{2}\right)^{1/2}/\sqrt{J(G+2)+G(G-1)/2}<\delta_{1},$
6:and $\max_{j\in\{1,...,J\}}h(|\lambda_{j2}^{(t)}|,...,|\lambda_{j,G+1}^{(t)}|)<\delta_{2}$ then
7:         Break
8:     end if
9:end while
10:$\Lambda^{(t)},\bm{\gamma}^{(t)},\bm{\psi}^{(t)}$.
The updating rule of $\beta^{(t)}_{jkk^{\prime}}$ and $c^{(t)}$ follows equation (1) and (47) in Chapter 2.2 of Bertsekas, (2014). The updating rule for $\beta_{jkk^{\prime}}^{(t)}$ follows the first-order optimality condition based on optimisations (4) and (5). The updating rule for $c^{(t)}$ ensures that it will become sufficiently large, which is necessary to guarantee the solution of the algorithm to converge to the feasible region defined by the zero constraints. On the other hand, it also prevents $c^{(t)}$ from growing too quickly with which the optimisation (5) is ill-conditioned. As shown in Chapter 2.2 of Bertsekas, (2014), as long as the sequences $\{\beta_{jkk^{\prime}}^{(t)}\}$ remain bounded, the sequence $\{c^{(t)}\}$ remains bounded. We follow the recommended choices of $c_{\theta}=0.25$ and $c_{\sigma}=10$ in Bertsekas, (2014), while pointing out that the performance of Algorithm 1 is quite robust against the choices of these tuning parameters; see Appendix C for a sensitivity analysis. The convergence of Algorithm 1 is guaranteed by Proposition 2.7 of Bertsekas, (2014).
@@
We stop the algorithm when both criteria are below their pre-specified thresholds, $\delta_{1}$ and $\delta_{2}$, respectively. The first criterion is a standard criterion for monitoring parameter convergence. This criterion being sufficiently small suggests the convergence of the algorithm. The second criterion is used to ensure that the solution is sufficiently close to the feasible set of optimisation defined by the equality constraints. This criterion being below $\delta_{2}$ means that for each variable $j$, there can only be one group loading whose absolute value is above the threshold $\delta_{2}$, and all the rest have absolute values below the threshold. Based on this, we can obtain an estimate of the bi-factor structure. More specifically, let $T$ be the last iteration number. Then the estimated bi-factor model structure is given by
@@
The optimisation problem (5) is non-convex and can get stuck in a local minimum. Thus, we recommend running the proposed algorithm multiple times with random starting points and choosing the solution with the smallest objective function value. The algorithm can also suffer from slow convergence, especially when the penalty term becomes large. When the algorithm does not converge within $T_{\max}$ iterations, we suggest using the estimated parameters at the $T_{\max}$th iteration as the initial parameters and restarting the optimisation until a good proportion of them converge. In the simulation study in Section 3 below, the estimated parameters obtained using 50 random starting points are close to the global minimum in most cases in the simulation study. For the real data example in Section 4, 200 random starting points are used to ensure a reliable result. We set $T_{\max}=100$ in all of our numerical studies.
@@ 3.1 Study I
In this study, we compare the proposed method with the oblique bi-factor rotation in Jennrich and Bentler, (2012) regarding their performance in the estimation accuracy of parameters and the recovery of the bi-factor structure. We consider two different settings for the data generation mechanism: (1) the observed data are generated from an exact bi-factor model, and (2) the observed data are generated from an approximate bi-factor model, where the loading matrix is generated by adding small perturbations to an exact bi-factor loading matrix.
The oblique bi-factor rotation method first estimates the loading matrix $\widehat{\Lambda}$ under the exploratory factor analysis setting by the optimisation problem
@@
We restrict $\lambda_{ij}=0$ for $i=2,\ldots,G$ and $j=i+1,\ldots,G+1$ to avoid the rotational indeterminacy of $\widehat{\Lambda}$, as suggested in Anderson and Rubin, (1956). Then the rotated solutions $\widehat{\Lambda}^{oblq}$ and $\widehat{\Phi}^{oblq}$ are obtained by finding a rotation matrix that solves the optimisation problem for oblique bi-factor rotation (Jennrich and Bentler,, 2012). The implementation in the R package GPArotation (Bernaards and Jennrich,, 2005) is used for solving this optimisation problem, which is based on a gradient projection algorithm. The optimisation problem for rotation is also nonconvex and thus may converge to local solutions. For a fair comparison, we also use 50 random starting points for the initial rotation matrix, which is the same as the number of random starting points that are used when running Algorithm 1.
We first examine the accuracy in estimating the loading matrix. We calculate the mean squared error (MSE) for $\widehat{\Lambda}$, after adjusting for the label and sign indeterminacy as considered in Theorem 1 and further discussed in Remark 1. More specifically, let $\mathcal{P}$ and $\mathcal{D}$ be the sets of permutation and sign flip matrices, respectively, as defined in Theorem 1. We define the MSE for $\widehat{\Lambda}$ as
$\text{MSE}_{\widehat{\Lambda}}=\min_{P\in\mathcal{P},D\in\mathcal{D}}\|\widehat{\Lambda}-\Lambda^{*}PD\|_{F}^{2}/(J(1+G)).$
Note that when data are generated from an approximate bi-factor model, $\Lambda^{*}$ does not have an exact bi-factor structure. This MSE is calculated for the loading matrix estimates from both methods.
@@
for $g=1,\ldots,G$ and some hard thresholding parameter $\delta>0$. In the analysis below, we consider three choices of hard thresholding parameter $\delta\in\{0.1,0.2,0.3\}$. We note that ^$\widehat{\mathcal{B}}_{g}^{oblq}$, $g=1,...,G$, may not yield an exact bi-factor structure as it is not guaranteed to return only one nonzero group loading parameter for each variable.
Let $\{\mathcal{B}_{g}^{*},g=1,\ldots,G\}$ be the true non-overlapping clusters of the $J$ variables, and let $\{\mathcal{B}_{g},g=1,\ldots,G\}$ be their estimates, either from the proposed method or the rotation method. When data are generated from an approximate bi-factor model, the true group clusters $\{\mathcal{B}_{g}^{*},g=1,\ldots,G\}$ are based on the corresponding bi-factor loading matrix before the perturbation. As the group factors can only be recovered up to label swapping, as Theorem 1 suggests, we measure the matching between the true and estimated structure up to a permutation of the factor labels. Specifically, the following two evaluation criteria are considered:
@@
  • Average Correctness Criterion (ACC): $\max_{\sigma\in\widetilde{\mathcal{P}}}\sum_{g=1}^{G}(|\mathcal{B}_{\sigma(g)}\cap\mathcal{B}_{g}^{*}|+|\mathcal{B}_{\sigma(g)}^{C}\cap\mathcal{B}_{g}^{*C}|)/(JG)$, which is the proportion of correctly identified zero and nonzero group loadings. Here for any set $\mathcal{B}$, $\mathcal{B}^{C}=\{1,\ldots,J\}\setminus\mathcal{B}$ is the complement of set $\mathcal{B}$.
Here, the EMC measures the perfect recovery of the true bi-factor structure, while the ACC can be viewed as a smooth version of EMC that measures the level of partial recovery. EMC = 1 when ACC = 1 and, EMC = 0 when ACC $<1$. A larger value of ACC indicates a higher level of partial recovery of the true bi-factor structure. More specifically, for a given permutation $\sigma\in\widetilde{\mathcal{P}}$, the quantity $|\mathcal{B}_{\sigma(g)}\cap\mathcal{B}_{g}^{*}|+|\mathcal{B}_{\sigma(g)}^{C}\cap\mathcal{B}_{g}^{*C}|$ computes the number of correctly identified nonzero and zero loadings for group factor $g$. For example, consider a case with $J=15$ items and $\mathcal{B}_{1}^{*}=\{1,4,7,10,13\}$ for the first group factor. If $\mathcal{B}_{\sigma(1)}=\mathcal{B}_{1}^{*}$, then for the first group factor, we have $|\mathcal{B}_{\sigma(1)}\cap\mathcal{B}_{1}^{*}|+|\mathcal{B}_{\sigma(1)}^{C}\cap\mathcal{B}_{1}^{*C}|=J=15$, i.e., all the nonzero and zero loadings have been correctly identified. If, instead, $\mathcal{B}_{\sigma(1)}=\{1,2,7,10,13\}$, then $|\mathcal{B}_{\sigma(1)}\cap\mathcal{B}_{1}^{*}|+|\mathcal{B}_{\sigma(1)}^{C}\cap\mathcal{B}_{1}^{*C}|=13$, i.e., 13 out of the 15 nonzero and zero loadings have been correctly identified. The quantity $\sum_{g=1}^{G}(|\mathcal{B}_{\sigma(g)}\cap\mathcal{B}_{g}^{*}|+|\mathcal{B}_{\sigma(g)}^{C}\cap\mathcal{B}_{g}^{*C}|)/(JG)$ thus computes the proportion of correctly identified zero and nonzero group loadings under the given permutation $\sigma$. The ACC considers all possible permutations of the group factor labels to account for the label indeterminacy.
To examine the recovery of the bi-factor structure, we consider $(J,G)\in\{(15,3),(30,5)\}$ and $N\in\{500,2000\}$. These choices, combined with the two settings for the data generation mechanism, lead to eight simulation settings. For each setting, we let $\mathcal{B}_{g}^{*}=\{g,g+G,\ldots,g+G(J/G-1)\}$ for $g=1,\ldots,G$, $\Psi^{*}=\mathbb{I}_{J\times J}$, and $\Phi^{*}=\Phi^{*}(\bm{\gamma}^{*})$ follow the reparameterization in Section 2.2, where the entries of $\bm{\gamma}^{*}$ are i.i.d., following a Uniform(0.5,0.5)0.50.5(-0.5,0.5)( - 0.5 , 0.5 ) distribution. Under the settings where data are generated from an exact bi-factor model, we generate the true loading matrix $\Lambda^{*}$ by
@@
for $j=1,\ldots,J$ and $k=1,\ldots,G+1$. In (9), $u_{jk}$s are i.i.d., following a Uniform$(0.2,1)$ distribution, and $x_{jk}$s are i.i.d., following a Bernoulli$(0.5)$ distribution. Under the settings where data are generated from an approximate bi-factor model, we generate the true loading matrix $\Lambda^{*}$ by
[…]
for $j=1,\ldots,J$ and $k=1,\ldots,G+1$. Here, $u_{jk}$s and $x_{jk}$s are generated in the same way as those in the exact bi-factor model, and $w_{jk}$ are i.i.d., following a Uniform$(0,0.1)$ distribution. In (10), the nonzero values of $\lambda_{jk}^{*}$ when $k>1$ and $j\notin\mathcal{B}_{k-1}^{*}$ represent the perturbation of $\Lambda^{*}$ from an exact bi-factor structure.
For each setting, we first generate $\Lambda^{*}$ and $\Phi^{*}$ once and use them to generate 100 datasets. The true parameter values for these simulations are given in Appendix B. The results about the estimation of the loading matrix are shown in Table 1(b). When data are generated from an exact bi-factor model, the proposed method outperforms the rotation method in terms of the MSE of the estimated loading matrix, as shown in Table 1(b)(a). When data are generated from an approximate bi-factor model, as shown in Table 1(b)(b), the proposed method is slightly better under the small-sample settings when $N=500$ but slightly worse under the large-sample settings when $N=2000$. The disadvantage of the proposed method under the large-sample settings is due to the bias brought by model misspecification. That is, the data generation model is not an exact bi-factor model, while the proposed method restricts its estimates in the space of exact bi-factor models.
The results about the recovery of the bi-factor structure based on the EMC and ACC metrics are shown in Tables 2(b) and 3(b), respectively. For the rotation method, the threshold $\delta=0.2$ yields the best results among the three threshold choices under all the simulation settings and for both performance metrics. However, even the results of the rotation method under this choice of threshold are not as good as those from the proposed method, especially when we look at the EMC metric. For example, when $J=30$, $G=5$, and $N=500$, the rotation method with $\delta=0.2$ can only correctly recover the entire bi-factor structure 15 times among 100 simulations, while the proposed method can correctly recover it 85 times.
Table 1: Simulation results of the MSE of $\widehat{\Lambda}$ estimated by the proposed ALM method and the exploratory bi-factor rotation method.
$(J,G)$ $N$ ALM Rotation
(15,3) 500 $2.10\times 10^{-3}$ $3.62\times 10^{-3}$
2000 $0.54\times 10^{-3}$ $0.92\times 10^{-3}$
(30,5) 500 $1.36\times 10^{-3}$ $4.94\times 10^{-3}$
2000 $0.30\times 10^{-3}$ $1.15\times 10^{-3}$
(a) The exact bi-factor model cases.
$(J,G)$ $N$ ALM Rotation
(15,3) 500 $4.74\times 10^{-3}$ $5.92\times 10^{-3}$
2000 $3.06\times 10^{-3}$ $2.40\times 10^{-3}$
(30,5) 500 $3.75\times 10^{-3}$ $3.88\times 10^{-3}$
2000 $2.63\times 10^{-3}$ $1.25\times 10^{-3}$
(b) The approximate bi-factor model cases.
Table 2: Simulation results of the EMC of the proposed ALM method and the exploratory bi-factor rotation method with three choices of hard thresholding parameter $\delta$.
$(J,G)$ $N$ ALM $\delta=0.1$ $\delta=0.2$ $\delta=0.3$
(15,3) 500 1.00 0.18 0.90 0.28
2000 1.00 0.99 1.00 0.50
(30,5) 500 0.85 0.00 0.15 0.00
2000 1.00 0.55 0.68 0.00
(a) The exact bi-factor model cases.
$(J,G)$ $N$ ALM $\delta=0.1$ $\delta=0.2$ $\delta=0.3$
(15,3) 500 0.99 0.00 0.62 0.33
2000 1.00 0.02 0.94 0.67
(30,5) 500 0.86 0.00 0.23 0.00
2000 1.00 0.00 0.82 0.00
(b) The approximate bi-factor model cases.
Table 3: Simulation results of the ACC of the proposed ALM method and the exploratory bi-factor rotation method with three choices of hard thresholding parameter $\delta$.
$(J,G)$ $N$ ALM $\delta=0.1$ $\delta=0.2$ $\delta=0.3$
(15,3) 500 1.000 0.966 0.998 0.976
2000 1.000 0.999 1.000 0.987
(30,5) 500 0.998 0.892 0.987 0.973
2000 1.000 0.996 0.998 0.980
(a) The exact bi-factor model cases.
$(J,G)$ $N$ ALM $\delta=0.1$ $\delta=0.2$ $\delta=0.3$
(15,3) 500 0.999 0.864 0.989 0.978
2000 1.000 0.928 0.999 0.991
(30,5) 500 0.998 0.848 0.989 0.978
2000 1.000 0.927 0.999 0.980
(b) The approximate bi-factor model cases.